OpenStudy (anonymous):

Someone Please Help!!!! Researchers count 30 tagged zebras and 150 total zebras in an area of a game reserve. If researchers had previously tagged 80 total zebras at the game reserve, about how many total zebras live at the game reserve? Write a proportion to estimate the total population.

OpenStudy (anonymous):

@campbell_st @thomaster Please help me with this......

OpenStudy (anonymous):

Try setting up your proportions as 30 over 150 equals 80 over x. Then, solve for x.

OpenStudy (anonymous):

Can you help me as in "What would x be?"

OpenStudy (anonymous):

To solve for x, you will need to cross-multiply. Are you familiar with that?

OpenStudy (anonymous):

no......Sorry im kinda new to this........

OpenStudy (anonymous):

ok, your equation when set up will look like this: \[\frac{ 30 }{ 150 }=\frac{ 80 }{ x }\]

OpenStudy (anonymous):

You will cross multiply to get: \[30x = 80*150\]

OpenStudy (anonymous):

do you know how to solve for x from there?

OpenStudy (anonymous):

Not really..... Can you help me a little bit more?

OpenStudy (anonymous):

Alright, multiply 80 and 150 together and tell me what you get.

OpenStudy (anonymous):

12000

OpenStudy (anonymous):

Alright, now how would you get X by itself?

OpenStudy (anonymous):

Do you divide12000 by 30?

OpenStudy (anonymous):

You sure do.

OpenStudy (anonymous):

And then the answer to that will be what x is?

OpenStudy (anonymous):

And that will give you your approximate number of total zebras, yes.

OpenStudy (anonymous):

Thank you so much!!!!

OpenStudy (anonymous):

No prob, take it easy.

OpenStudy (anonymous):

Ok, you 2